OpenStudy (anonymous):

what is the tangent line of function f(x)=x^3-6x^2+11x-6. the derivative of f(x)is f'(x)=3x^2-12x+11

OpenStudy (anonymous):

f(4)=6 so the tangent line in question has equation y-6=11(x-4), or y=11x-38. Which one is it?

OpenStudy (anonymous):

can anyone help me out?